# LVBench: An Extreme Long Video Understanding Benchmark

### v2

@@ Abstract
Figure 1: (Left) Video categories. Our dataset contains 6 major categories and 21 subcategories. (Right) Performance radar chart of different models on LVBench.
@@ 1 Introduction
Recently, the rapid advancements in large language models (OpenAI, 2023; Anthropic, 2024; Du et al., 2021) and visual feature extraction models (Radford et al., 2021; Sun et al., 2023; Zhai et al., 2023) have led to significant improvements in the performance of multimodal large models on open-domain video question-answering tasks. These multimodal understanding models have also empowered various downstream tasks, such as embodied intelligence, video generation, and subtitles for the visually impaired. However, most current end-to-end video understanding models are limited to processing videos of only a few minutes in length. More complex tasks require the capability to understand much longer videos, which presents a significant challenge to existing multimodal models.
Despite numerous video understanding benchmarks being proposed in the past, the field of long video understanding remains underdeveloped due to the difficulties in data acquisition and annotation. To address this gap, we introduce LVBench, a benchmark designed to evaluate the capabilities of models in understanding long videos. We collected a substantial amount of long video data from public sources, with annotations provided through a combination of manual effort and model assistance. Additionally, we carefully designed a series of evaluation tasks. Compared to previous video understanding benchmarks (Li et al., 2023c), LVBench offers the following unique features:
  • We define six core capabilities for long video understanding, which can be flexibly combined to create complex and challenging questions. This multifaceted approach enables a comprehensive evaluation of a model’s ability to process and comprehend lengthy video content.
  • We have collected a diverse range of long video data from various sources, with an average duration approximately four times longer than the longest existing dataset. The categories of videos in LVBench are illustrated in Figure 1. This extensive collection of long-form video content provides a robust foundation for testing models on extended temporal contexts.
  • Through meticulous human annotation and multi-stage quality control processes, we ensure the high quality of our dataset, providing a reliable benchmark for assessing long video understanding capabilities.
@@ 2 Related Works
Multi-modal Large Language Models. Building upon the achievements in Large Language Models (LLMs), the field has shifted towards Multi-modal Large Language Models (MLLMs) to enhance multi-modal understanding and generation capabilities (Wang et al., 2023; Hong et al., 2023; Alayrac et al., 2022; Li et al., 2023b; c; Liu et al., 2024c; Xu et al., 2024). Early advancements in this area include models like Flamingo (Alayrac et al., 2022), which fused text and vision to perform exceptionally well in multimodal tasks. Subsequent models such as VideoChat (Li et al., 2023b) and VideoChatGPT (Maaz et al., 2024) began exploring the video modality, using ChatGPT (Achiam et al., 2023) to generate video instruction-tuning data for improved instruction-following capabilities. VideoChat2 (Li et al., 2023c) advanced the field by introducing a dedicated video encoder, requiring extensive training on large-scale datasets to optimize performance. The ST-LLM (Liu et al., 2024c) model streamlined this process by leveraging LLMs for visual sequence modeling, thereby reducing training complexity and enhancing performance. PLLaVA (Xu et al., 2024) introduced a resource-efficient method for adapting image-language pre-trained models to dense video understanding through a novel feature pooling strategy, achieving state-of-the-art results. Gemini 1.5 Pro (Reid et al., 2024) further pushed the boundaries with a mixture-of-experts architecture, excelling in long-context reasoning and multi-modality across extensive multimodal benchmarks. These advancements underscore the significant progress and potential of MLLMs in advancing multimodal comprehension and generation. Despite the progress made, our experiments indicate that current video understanding models still fall short on tasks requiring long-range comprehension, highlighting an urgent need for the development of models tailored for long video understanding.
Benchmarks for MLLM. Recent advancements in vision-language (VL) benchmarks have largely focused on images and short videos, as seen in datasets like MMBench (Liu et al., 2023), SEED-Bench-2 (Li et al., 2023a), TGIF-QA (Jang et al., 2017) and MVBench (Li et al., 2023c). For long video understanding, previous benchmarks such as Perception Test (Pătrăucean et al., 2023) have explored multi-modal video perception and reasoning but often with shorter video clips and limited temporal context. Datasets like How2QA (Li et al., 2020) and ActivityNet-QA (Yu et al., 2019) are domain-specific and do not adequately capture the complexity of long-term video understanding. EgoSchema (Mangalam et al., 2024) and MovieQA (Tapaswi et al., 2016) provide insights into narrative and thematic understanding but are constrained by shorter video durations and limited granularity. While LongVideoBench (Wu et al., 2024), MovieChat (Song et al., 2023), MoVQA (Zhang et al., 2023), and Video-MME (Fu et al., 2024) utilize longer videos to test models, their average duration is still limited to around 10 minutes. In contrast, LVBench features significantly longer video segments averaging 4101 seconds, pushing the boundaries of long-term video understanding with comprehensive tasks and detailed annotations.
@@
[…]
@@ 3.1 Dataset Collection
We define a long video as having the following properties:
  • A duration exceeding 30 minutes.
  • Highly dynamic content with rich visual information.
To curate our dataset, we sourced publicly available videos from YouTube, covering a diverse range of topics such as sports, live streams, TV shows, documentaries, animations, and more. By using various search terms and YouTube’s auto-suggestion feature, we gathered an initial collection of 500 videos, each with a minimum duration of 30 minutes. Subsequently, our annotators carefully screened these videos based on the following criteria to select a subset of 103 high-quality, diverse videos:
  • The presence of one or more protagonists (possibly in a first-person perspective) who serve as narrators, appearing on-screen for a significant portion of the video and interacting with the environment.
  • A complete video structure with a coherent logical flow.
  • The occurrence of multiple minor events throughout the video, following a chronological order and exhibiting completeness.
  • Visuals that are relatively easy to comprehend without overly fragmented information.
  • Video content that can be understood independently of audio cues.
This multi-stage filtering process ensures that our dataset comprises diverse, high-quality long videos suitable for evaluating complex video understanding tasks.
3.2 Task Definition
To comprehensively evaluate long video understanding, we propose a benchmark testing six core capabilities. Example questions for each capability are presented in Figure 2. The proportion of each capability is shown in Figure 3. Questions are designed to flexibly combine multiple skills to construct complex queries that probe a model’s capacity to:
  1. 1.
@@ 3.2 Task Definition
    Temporal Grounding(TG): Questions focus on understanding sequences and dynamics within the video, such as identifying specific events at designated times (e.g., "What happened at 29:30?").
  2. 2.
    Summarization(Sum): Annotators are required to produce an abstractive summary that encapsulates the entire video content, demonstrating a cohesive understanding of the sequence from start to finish.
  3. 3.
    Reasoning(Rea): This involves the application of advanced reasoning skills to interpret the video content:
    • Causal: Determining causal relationships between events.
    • Emotional: Understanding the emotional developments of characters.
    • Intentional: Interpreting the underlying intentions of characters.
    • Prospective: Making predictions about future events based on current evidence.
  4. 4.
    Entity Recognition(ER): This capability requires the identification and continuous tracking of key entities (such as people, places, and objects) throughout the video:
    • Entity Detection: Identifying mentions of entities and resolving their identities across different instances.
    • Relation Extraction: Extracting the relationships among identified entities.
    • Action Recognition: Observing and understanding the progression of an entity’s actions over time.
    • Entity Association: Linking entities to relevant events.
  5. 5.
    Event Understanding(EU): Comprehending overarching semantic concepts in the video:
    • Video Classification: Classifying the genre of the video (e.g., news, film).
    • Event Detection: Identifying significant occurrences within the video (such as goals scored or conflicts).
    • Scene Perception: Recognizing changes between different scenes or settings.
  6. 6.
@@
    Key Information Retrieval(KIR): Extracting specific, detailed information, such as text displayed in the video (e.g., "What revenue growth did the firm report at the conference?").
By composing questions that test combinations of these temporal, abstractive, reasoning, entity-centric, event-based, and detail-oriented skills, our benchmark enables robust evaluation of a model’s ability to understand long videos across multiple modalities. This multifaceted typology covers the key cognitive capabilities required to comprehend complex, open-ended video.
@@
3.3 QA Generation
The total number of questions for each video is positively correlated with the video duration, averaging 24 questions per hour. After constructing a question, annotators are asked to provide four options, including one correct answer and three distractors. Annotating long videos is significantly more challenging than annotating short videos or image data, making quality control a substantial challenge. To ensure the quality of the evaluation set, we encourage annotators to follow these principles:
  • Question Diversity: Construct at least one question for each question type in a video.
  • Specificity: Questions should refer to unique scenes, events, or characters, avoiding vague descriptions. For example, if a video contains two arguments, a well-constructed question might be: "When did A and B start arguing?" or "How did the person in red’s expression change during the hallway argument?". Less specific questions would be: "Why did they start arguing?" or "Who are the people arguing?"
  • Temporal Coverage: Questions should cover multiple events throughout the video, avoiding repetition of a single event.
  • Consistency: Constructed correct answers should precisely address the questions. Answers should match the content of the questions, avoiding irrelevant information. Correct and incorrect answers should be constructed consistently, avoiding obvious differences in length, form, or format.
By following these principles, we aim to create a high-quality, diverse, and challenging evaluation set for long video understanding.
@@ 3.4 Data Quality Control
Upon constructing all the questions, we discovered that for certain questions, a language model could generate answers without any visual input. As highlighted in MMstar (Chen et al., 2024a), many multimodal benchmarks can be effectively solved using pure text input alone. To mitigate this issue, we employed a straightforward yet effective approach. We utilized two powerful large language models, GLM-4 (Du et al., 2021) and GPT-4 (Achiam et al., 2023), to independently generate answers for all the questions. In cases where the outputs from both models were identical and matched the ground truth answer, we removed that particular data sample from the dataset. This filtering process successfully eliminated the majority of questions that did not rely on video content for answering. Following this filtering step, we obtained a refined set of 1,549 question-answer pairs.

### v3

@@ 1 Introduction
The rapid development of Multi-modal Large Language Models (MLLMs) has led to significant advances in vision-language understanding [26, 37, 38], driven by breakthroughs in both large language models [25, 3, 8] and visual encoders [30, 35, 46]. These models have demonstrated impressive capabilities across various video understanding tasks, from fundamental tasks like action recognition and object tracking, to more complex tasks such as temporal event localization, dense video captioning, and abstractive video summarization. However, while existing MLLMs excel at processing short video clips, they face substantial challenges when dealing with longer temporal sequences - a crucial limitation that hinders their practical applications. Despite numerous video understanding benchmarks being proposed, most primarily focus on evaluating spatial understanding in static images or short video clips, overlooking the critical aspect of temporal understanding in long-form videos. This gap is particularly significant given that real-world applications often require comprehension of extended temporal contexts. The scarcity of long video understanding benchmarks can be attributed to the challenges in data collection and the complexity of annotation. To address these limitations, we introduce LVBench, a comprehensive benchmark designed specifically for evaluating MLLMs’ capabilities in long video understanding. Our benchmark distinguishes itself from existing datasets through several key features:
  • We establish a systematic framework of six core temporal understanding capabilities, which can be flexibly combined to create complex, multi-faceted evaluation tasks. This design enables a thorough assessment of models’ ability to process and reason about extended temporal sequences.
  • We curate a diverse collection of long-form videos from various domains, with an average duration approximately four times longer than existing benchmarks. The videos span multiple categories (as illustrated in Figure 1), providing a robust foundation for evaluating temporal understanding across different contexts.
  • Through a rigorous combination of expert human annotation and quality control processes, we ensure high-quality ground truth annotations, establishing a reliable benchmark for assessing long video understanding capabilities.
@@ 2 Related Works
TGIF-QA [14] 165,165 3 $usym{2713} \usym$$\usym{2717}$ Auto
MSVD-QA [42] 13,157 10 $usym{2713} \usym$$\usym{2717}$ Auto
MSRVTT-QA [42] 72,821 15 $usym{2713} \usym$$\usym{2717}$ Auto
MVBench [17] 4,000 16 $usym{2713} \usym$$\usym{2713}$ Auto
@@
NExT-QA [41] 52,044 44 $usym{2713} \usym$$\usym{2717}$ Manual
How2QA [18] 44,007 60 $usym{2713} \usym$$\usym{2713}$ Manual
@@
CinePile [31] 303,828 160 $usym{2717} \usym$$\usym{2713}$ Auto&Manual
EgoSchema [24] 5,000 180 $usym{2717} \usym$$\usym{2713}$ Auto&Manual
MovieQA [36] 6,462 203 $usym{2717} \usym$$\usym{2713}$ Manual
LongVideoBench [40] 6,678 473 $usym{2713} \usym$$\usym{2713}$ Manual
MovieChat-1K [34] 13,000 564 $usym{2717} \usym$$\usym{2713}$ Manual
MoVQA [49] 21,953 992 $usym{2717} \usym$$\usym{2713}$ Manual
Video-MME [9] 2,700 1018 $usym{2713} \usym$$\usym{2713}$ Manual
LVBench(Ours) 1,549 4,101 $usym{2713} \usym$$\usym{2713}$ Manual
2.1 Multi-modal Large Language Models
Building upon the achievements in LLMs, the field has shifted towards MLLMs to enhance multi-modal understanding capabilities [2, 11, 38, 26, 32]. Early advancements in this area include models like Flamingo [2], which introduced a novel approach by adding cross-attention layers to a frozen language model, enabling efficient vision-language alignment without full model retraining. Recent works have made significant progress in video understanding through various architectural innovations. Video-LLaMA [48] employs a video transformer with Q-Former for effective temporal modeling. Models like InternVL2 [6], LLaVA-Next [50], and CogVLM2 [12] unify high-resolution images and videos as multi-frame sequences, enabling seamless processing of both modalities. Qwen2-VL [37] introduces 3D-RoPE to achieve strong temporal extrapolation capabilities for long video understanding. To handle extended temporal contexts, MovieChat [34] proposes a dual-memory architecture with short-term and long-term memory modules that can efficiently process videos over 10,000 frames. LWM[20] with ring attention has pushed context lengths to the million-token scale, providing a promising foundation for long video understanding. PLLaVA[43] introduced a resource-efficient method for adapting image-language pre-trained models to dense video understanding through a novel feature pooling strategy. Despite these significant architectural advances, our experiments indicate that current video understanding models still fall short on tasks requiring long-range comprehension, highlighting an urgent need for the development of models specifically tailored for long video understanding.
2.2 Benchmarks for MLLM
Recent advancements in vision-language benchmarks have largely focused on images and short videos, as seen in datasets like MMBench [22], SEED-Bench-2 [15], TGIF-QA [14] and MVBench [17]. For long video understanding, previous benchmarks such as Perception Test [29] have explored multi-modal video perception and reasoning but often with shorter video clips and limited temporal context. Datasets like How2QA [18] and ActivityNet-QA [45] are domain-specific and do not adequately capture the complexity of long-term video understanding. EgoSchema [24] and MovieQA [36] provide insights into narrative and thematic understanding but are constrained by shorter video durations and limited granularity. While LongVideoBench [40], MovieChat [34], MoVQA [49], and Video-MME [9] utilize longer videos to test models, their average duration is still limited to around 10 minutes. In contrast, LVBench features significantly longer video segments averaging 4101 seconds, pushing the boundaries of long-term video understanding with comprehensive tasks and detailed annotations. A detailed comparison between LVBench and existing video benchmarks is presented in Table 1.
@@ 3.1 Dataset Collection
We define long videos as those having a minimum duration of 30 minutes and containing rich, dynamic visual information with multiple events and scene transitions. To construct a comprehensive benchmark for long video understanding, we curated a diverse collection of videos from YouTube through the following systematic process.
Video Collection. Starting with YouTube’s search and recommendation algorithms, we gathered an initial pool of 500 videos using carefully selected keywords across different domains. Our search terms covered six primary categories: Sports Competitions (e.g., basketball matches and Olympic games), Documentary Films (e.g., nature documentaries and historical records), Event Records (e.g., ceremonies and festivals), Lifestyle and Daily Activities (e.g., cooking tutorials and travel vlogs), TV Shows and Drama Series (e.g., sitcoms and reality shows), and Cartoon Videos (e.g., animated movies and cartoon series).
Quality Filtering. Our annotators carefully screened these videos using five well-defined criteria:
  • Clear Protagonist Presence: Videos must feature identifiable main characters (human or virtual) who appear consistently and interact meaningfully with their environment. For first-person perspective videos, the narrator’s actions and decisions should be clearly observable.
  • Structural Coherence: Videos should maintain a well-defined narrative structure with clear beginning, development, and conclusion phases. We define “coherent logical flow” as the presence of causally connected events that progress naturally through time.
  • Event Density: Videos must contain multiple distinguishable events (at least one significant event every 5 minutes on average) arranged in chronological order. We define events as meaningful actions, interactions, or state changes that contribute to the video’s narrative.
  • Visual Clarity: The visual content should be of high quality (minimum 720p resolution) with stable camera work and clear scene compositions. We exclude videos with excessive rapid cuts, shaky footage, or poor lighting conditions.
  • Modality Independence: While audio may enhance understanding, all critical information should be conveyed through visual channels to ensure fair evaluation of vision-language models.
After applying these filtering criteria, we retained 103 high-quality videos totaling 117 hours of content.
3.2 Task Definition and Question Types
To systematically evaluate models’ capabilities in understanding long videos, we define a comprehensive taxonomy of six core skills essential for video comprehension. Each skill category is carefully designed to test specific aspects of temporal and semantic understanding while allowing for compositional combinations that assess more complex reasoning abilities.
@@ 3.2 Task Definition and Question Types
Temporal Grounding (TG). These questions assess the model’s ability to locate and understand events within the video’s timeline, such as identifying specific moments (“What happened at 29:30?”), determining event durations, and recognizing event sequences. This capability is fundamental for understanding the temporal dynamics in long videos.
Summarization (Sum). Tests the ability to synthesize and abstract information across the entire video content. Questions in this category require models to generate concise overviews, identify key developments, and understand the video’s central themes, demonstrating a comprehensive understanding of long-form content.
Reasoning (Rea). Evaluates higher-order cognitive skills including causal understanding (“Why did the experiment fail?”), emotional comprehension, intention interpretation, and future prediction. This category specifically examines a model’s ability to make complex inferences beyond surface-level observations.
Entity Recognition (ER). Focuses on identifying and tracking key entities (people, objects, places) throughout the video. This includes recognizing distinct entities, understanding their relationships, tracking their actions, and connecting them to relevant events - skills crucial for following long-form narratives.
Event Understanding (EU). Tests comprehension of high-level semantic concepts by requiring models to classify video types, detect significant events and understand scene transitions. This capability is essential for grasping the overall structure and flow of long videos.
@@
Key Information Retrieval (KIR). Evaluates attention to specific details by requiring extraction and comprehension of visual text, numerical data, and other precise information presented in the video, such as “What revenue growth did the firm report at the conference?”
@@
Figure 2 provides the distribution of each question type. As shown in the figure, Entity Recognition (ER) and Event Understanding (EU) account for a larger proportion of the questions. This imbalance naturally arises from the inherent structure of video content, where almost all questions inevitably involve identifying or tracking entities and understanding events or actions. Consider the question “Why did Player number 4 in white push down Player number 17 in purple during the match?”. While this question primarily tests reasoning capabilities (understanding the motivation behind an action), it inherently requires entity recognition (identifying and distinguishing the two players by their numbers and jersey colors) and event understanding (recognizing the pushing action within the context of the match).
3.3 Question-Answer Pair Generation
The annotation of long videos presents unique challenges compared to short-form content, requiring careful consideration of temporal dynamics and comprehensive coverage. We developed a systematic annotation protocol consisting of three key stages to ensure high-quality, diverse, and challenging question-answer pairs.
Video Analysis. In Stage 1 (Video Analysis), annotators first watch each video in its entirety. During this initial viewing, they mark significant events, scene transitions, and key entities, while also identifying potential temporal dependencies and causal relationships that could form the basis for meaningful questions.
Question Generation. Stage 2 (Question Generation) focuses on creating questions that thoroughly probe the video content. The number of questions scales with video duration, averaging 24 questions per hour. Questions are strategically distributed across the entire video duration to ensure comprehensive temporal coverage. Each video must include questions testing all six core capabilities. In constructing these questions, we emphasize specificity to ensure unique and unambiguous references to video content. For example, when questioning an argument scene, we prefer precise formulations like “When did A and B start arguing?” or “How did the person in red’s expression change during the hallway argument?” over vague alternatives such as “Why did they start arguing?” or “Who are the people arguing?”. This specificity principle ensures that questions have clear, uniquely identifiable answers within the video context.
Answer Construction. Stage 3 (Answer Construction) focuses on both creating answer choices and annotating temporal information needed for answering each question. For each question, annotators create four answer choices - one correct answer and three distractors. The correct answer precisely addresses the question, while distractors are carefully designed to share similar format and length with the correct answer. These distractors incorporate plausible but incorrect information from the video and test different aspects of understanding.
Importantly, during this stage, annotators also manually identify and mark what we term the “clue duration” - the minimum video segment needed to accurately answer each question. This temporal annotation involves carefully reviewing the video content to determine the exact start and end timestamps of the critical segments containing the information required for answering the question. For instance, a question about causal relationships might require a longer clue duration spanning multiple related events, while a question about a specific visual detail might have a shorter clue duration focused on a single scene.
Table 2: LVBench evaluation results regarding each core long video understanding capability. The highest score is highlighted with green, and the second highest is highlighted with purple. All the numbers are presented in % and the full score is 100%.
Model LLM ER EU KIR TG Rea Sum Overall
Non-Native Long Video Support Models
TimeChat [33] LLaMA2-7B 21.9 21.7 25.9 22.7 25.0 24.1 22.3
Video-ChatGPT [23] Vicuna-1.5-13B 22.9 22.6 22.7 25.5 23.4 24.1 23.1
PLLaVA [43] Yi-34B 25.0 24.9 26.2 21.4 30.0 25.9 26.1
LLaVA-OneVision [16] LLaMA3-70B 25.0 26.9 29.2 30.9 25.4 31.0 26.9
CogVLM2-Video [12] LLaMA3-8B 28.3 26.9 31.0 25.1 25.5 38.9 28.1
LLaVA-NeXT [50] Yi-34B 30.1 31.2 34.1 31.4 35.0 27.6 32.2
InternVL2-40B [6] Nous-Hermes-2-Yi-34B 37.4 39.7 43.4 31.4 42.5 41.4 39.6
mPLUG-Owl3 [44] Qwen2-7B 46.0 41.6 42.4 41.1 47.5 40.4 43.5
GLM-4.1V-Thinking [13] GLM-4-9B 42.8 43.5 50.0 39.1 46.0 25.9 44.3
VideoLLaMA3-7B [47] Qwen2.5-7B 45.8 42.4 47.8 35.9 45.8 36.2 45.3
GLM4V-Plus [12] GLM-4 46.2 47.8 54.1 42.7 46.5 37.9 48.7
GPT-4o-20241120 [26] GPT-4o 48.9 49.5 48.1 40.9 50.3 50.0 48.9
GPT-4.1 [27] GPT-4.1 - - - - - - 60.1
Native Long Video Support Models
MovieChat [34] Vicuna-7B 21.3 23.1 25.9 22.3 24.0 17.2 22.5
LLaMA-VID [19] Vicuna-13B 25.4 21.7 23.4 26.4 26.5 17.2 23.9
LWM [20] LLaMA2-7B 24.7 24.8 26.5 28.6 30.5 22.4 25.5
Gemini-1.5-Pro [32] Gemini 1.5 Pro 32.1 30.9 39.3 31.8 27.0 32.8 33.1
Kangaroo [21] LLaMA3-8B 38.6 37.9 29.6 35.0 41.3 36.2 38.3
Qwen2-VL-72B [37] Qwen2-72B 38.0 41.1 38.3 41.4 46.5 46.6 41.3
Qwen2.5-VL-72B [4] Qwen2.5-72B 44.2 40.9 55.6 37.7 45.2 34.5 44.0
Gemini-2.0-Flash [7] Gemini-2.0-Flash 47.4 48.5 56.8 39.3 44.4 41.4 48.6
AdaReTaKe [39] Qwen2.5-72B 53.0 50.7 62.2 45.5 54.7 37.9 53.3
Gemini-2.5-Flash [7] Gemini-2.5-Flash 55.2 55.5 63.8 52.7 55.5 44.8 56.7
MR.Video [28] Gemini-2.0-Flash 59.8 57.4 71.4 58.8 57.7 50.0 60.8
Seed1.5-VL [10] Seed1.5 64.3 64.0 64.7 52.3 65.0 51.7 64.0
Seed1.5-VL-Thinking [10] Seed1.5 65.4 63.4 68.0 53.6 63.7 46.6 64.6
Gemini-2.5-Pro [7] Gemini-2.5-Pro 64.5 67.5 72.8 65.9 66.5 58.6 67.4
@@ 3.4 Data Quality Control
Upon constructing all the questions, we discovered that for certain questions, a language model could generate answers without any visual input. As highlighted in MMstar [5], many multimodal benchmarks can be effectively solved using pure text input alone. To mitigate this issue, we employed a straightforward yet effective approach. We utilized two powerful large language models, GLM-4 [8] and GPT-4 [1], to independently generate answers for all the questions. In cases where the outputs from both models were identical and matched the ground truth answer, we removed that particular data sample from the dataset. This filtering process successfully eliminated the majority of questions that did not rely on video content for answering. Following this filtering step, we obtained a refined set of 1,549 question-answer pairs.
Figure 3: Failure analysis of native long-video models via answer distribution.